# Physics-Constrained Fine-Tuning of Flow-Matching Models for Generation and Inverse Problems

### v1

@@ Appendix D Details: Experimental Results
Runtime $\approx$12 h
Here the fine-tuning network has $\approx 39$ M parameters, the inverse predictor remains at $\approx 9$ M. In 3D, especially when upsampling by a factor of 4, test function evaluation and network evaluations are significantly more expensive than the respective operations in 2D.
Point-wise Distribution Metrics
Following Kerrigan, Migliorini, and Smyth (2023), we compare data and model by averaging statistics which describe similarity at every grid point. Denote by $\mu_{\mathrm{data}}(\xi),\;\sigma_{\mathrm{data}}(\xi)$ the empirical mean and standard deviation at location $\xi\in\Omega$ and by $\mu_{\mathrm{gen}}(\xi),\;\sigma_{\mathrm{gen}}(\xi)$ the corresponding quantities estimated from generated samples. All scores are normalised by the average data variance
$S_{\mathrm{data}}=\tfrac{1}{|\Omega|}\sum_{\xi}\sigma_{\mathrm{data}}^{2}(\xi).$
Mean mismatch
$\mathrm{MMSE}_{\mathrm{rel}}=\frac{1}{S_{\mathrm{data}}|\Omega|}\sum_{\xi}\bigl{(}\mu_{\mathrm{gen}}-\mu_{\mathrm{data}}\bigr{)}^{2}.$
Spread mismatch
$\mathrm{SMSE}_{\mathrm{rel}}=\frac{1}{S_{\mathrm{data}}|\Omega|}\sum_{\xi}\bigl{(}\sigma_{\mathrm{gen}}-\sigma_{\mathrm{data}}\bigr{)}^{2}.$
Ensemble diversity
$\mathrm{Diversity}_{\mathrm{rel}}=\frac{1}{S_{\mathrm{data}}|\Omega|}\sum_{\xi}\sigma_{\mathrm{gen}}^{2}(\xi).$
For the ensemble diversity, a score of $1$ indicates parity with the data; values <1<1 denote less variability, whereas values >1>1 reveal increased point-wise variance in generated samples. All three metrics are dimensionless and therefore comparable across datasets and resolutions.
Appendix E Additional Results
Denoising
Details on Denoising with $\lambda_{f}=0$.
To further illustrate the effectiveness of our proposed regularization, we here demonstrate that similar results cannot be obtained by simply reducing the fine-tuning strength $\lambda_{x}$. Figure 12 shows fine-tuning results for $\lambda_{f}=0$ and $\lambda_{x}\in\{100,1000,5000\}$. For small $\lambda_{x}$, fine-tuning retains some similarity to the sample of the base model, but does not lead to fully denoised samples. On the other hand, fully denoised samples for high values of $\lambda_{x}$ have diverged far from the samples of the base model, and, as Table 2 indicates, show signs of mode collapse.
Figure 12: Fine-tuned models for pre-trained model on noisy data without regularization, $\lambda_{f}=0$ for different fine-tuning strengths $\lambda_{x}$. The top row shows the pressure distribution induced by the permeability in the second row.
Regularization Strength $\lambda_{f}$.
For a fixed $\lambda_{x}=15\text{k}$, we here compare results from fine-tuning with different regularization strengths $\lambda_{f}\in\{0.001,0.01,0.1\}$. Here, all models are able to produce visually denoised samples, see Fig. 13. However, for lower values of regularization we again see a higher deviation from the base sample, both for the obtained pressure and permeability, similar to higher $\lambda_{x}$ in the previous example.
Figure 13: Fine-tuned models for pre-trained model on noisy data with different levels of regularization $\lambda_{f}$ for the same $\lambda_{x}$. The top row shows the pressure distribution induced by the permeability in the second row.
Residual Heatmaps.
We can visually represent the weak residuals as heatmaps by computing the inner product $\langle\mathcal{L}_{a}u,\psi^{(i)}\rangle_{L^{2}(\Omega)}$ for each spatial location with a test function centered at the respective coordinates. In this case, to ensure comparability, length-scales are not sampled but fixed. Figure 14 shows such heatmaps the fine-tuned model trained on the denoising task. Note that the model was fine-tuned with $\sigma\in[3,\,10]$, but the residual was also decreased for larger length-scales.
Figure 14: Heatmaps of residual for fine-tuned model on denoising task, showing residuals computed with two distinct length-scales for the test functions.
Guidance
We provide more samples obtained from guided samples based on sparese observations on the parmeter. Here, we show samples for $m\in\{10,100,1000\}$ in, Fig 19, Fig 20, and Fig 21 respectively. As expected, the variability of generated samples reduces drastically with an increased number of observations. Sparse observations are sampled from ground-truth training data. As indicated in Table 2, the base Flow Matching models show a reduced variability in the recovered parameter - an issue which might be resolved by for example using more training data or stronger base models. Therefore, we cannot expect that the model can be guided perfectly towards all parameter fields in the training data. However, results from guided generation follow the sparse observations quite accurately.
Acoustics
Figure 15 shows a sample from the base FM model (with speed of sound recovered from pre-trained inverse predictor) and the fine-tuned model on four times upsampled data.
Figure 15: Top row: generated sample from base FM model trained on acoustic wave data of resolution $64\times 64\times 64$. Second row: sample from fine-tuned model on upsampled data ($256\times 256\times 256$) for same initial noise. Time indices for sample from base model are adapted to match fine-tuned ones.
Since the scalar boundary loss reported in the main paper aggregates violations of the reflective condition over space and time, localised spikes can be obscured by averaging. To reveal the full spatio-temporal behaviour we visualise, in Fig. 16, the evolution of the squared normal-gradient magnitude $|\nabla_{\!n}p|^{2}$.
For a generated trajectory we concatenate the four spatial edges of the upsampled output into a single “boundary index” axis of length 4(H2)=10164\,(H-2)=1016 and stack the resulting vectors over the 256 time steps, producing a $256\times 1016$ matrix per model. We display the residuals for the base and fine-tuned model and the difference (base - fine-tuned). It can be seen that the fine-tuned model effectively suppresses boundary spikes uniformly across time.
[…]

### v2

@@ E.2 Ablations and Baselines
PBFM.
As an alternative baseline we adapt Physics-Based Flow Matching (PBFM) (Baldan et al., 2025) to our setting. Since PBFM requires paired $(x,\alpha)$ information during training, we augment the method with our pretrained inverse predictor $\varphi$ to estimate $\alpha$ and enable residual computation. Training follows the original ConFIG (Liu et al., 2024) update rule: at each iteration we compute the FM loss and the residual loss, extract their gradients, and form a conflict-free update via ConFIG. Our implementation mirrors the official code except for minor modifications to support complex-valued weights (arising from FNO backbones). Due to memory constraints, we have to use much lower batch sizes than with vanilla Flow Matching (16 instead of 128). We chose the number of epochs such that the total number of update steps is twice that of our base FM models.
PBFM serves as a comparison against a training-time physics-regularized baseline, whereas our method enforces physics post-training (data-free). Note that all our PDE settings deliberately introduce a mismatch between the physics assumed during fine-tuning and the physics underlying the training data. Such misspecification is inherently challenging for training-time methods like PBFM which naturally places them at a disadvantage.
E.3 Darcy
For the Darcy experiments, we use the hyperparameters listed in Table 7. Note that one epoch amounts to exactly one gradient descent step, meaning that we only fine-tune for 20 total steps. The parameters $\lambda_{x},\lambda_{\alpha}$ and $\lambda_{f}$ are varied in the experiments, see Figure 3 in the main text.
Table 7: Darcy fine-tuning hyperparameters.
Hyperparameter Value
Time–tilting factor $q$ 0.9
$\lambda_{x}$ varying
$\lambda_{\alpha}$ varying
$\lambda_{f}$ varying
$K_{\text{last}}$ 20
$K$ 20
Noise scaling $\kappa$ 0.9
Sampling steps (per trajectory) 100
Training epochs 20
Learning rate 0.00002
Batch size 15
E.4 Guidance
For generating the guidance results, we use the same model as in the Darcy experiments to further highlight that we can infer functional joint distributions when starting from noisy observations. Instead of the usual Euler stepping, here we use a Heun sampler following Huang et al. (2024). While this is more expensive, since we need to differentiate through two forward passes of our model to obtain the guidance gradient, it empirically improved faithfulness to the sparse observations significantly. Note, however, that we only guide on sparse observations and not on PDE residuals. The full sampling procedure is show in Algorithm 2. In our experiments, we use 100 steps for sampling and guidance scales of $\gamma_{x}=\gamma_{\alpha}=0.75.$
Algorithm 2 Guided Heun sampler with sparse observations
1:initial state $x_{0}$; initial parameter $\alpha_{0}$; observed targets $\alpha_{\mathrm{obs}}$; index set $\mathcal{I}$; steps $S$; guidance scales $\gamma_{x},\gamma_{\alpha}$; base fields $v^{x}_{t},\,v^{\alpha}_{t}$; fine-tuned joint field $v^{\text{ft}}_{t}$
2:trajectories $\{x^{(i)}\}_{i=0}^{S}$ and $\{\alpha^{(i)}\}_{i=0}^{S}$
[…]
7:
8:for $i=0$ to S1S-1 do
9:  Predictor (Euler):
10:  $v^{x}\leftarrow v^{x}_{t_{i}}\!\bigl(x^{(i)}\bigr)$
11:  $v^{\alpha}_{\mathrm{base}}\leftarrow v^{\alpha}_{t_{i}}\!\bigl(x^{(i)},\,\alpha^{(i)},\,v^{x}\bigr)$
12:  $(\tilde{v}^{x},\,\tilde{v}^{\alpha})\leftarrow v^{\text{ft}}_{t}\bigl(x^{(i)},\,\alpha^{(i)},\,v^{\alpha}_{\mathrm{base}}\bigr)$
13:  $\hat{x}\leftarrow x^{(i)}+h\,\tilde{v}^{x}$
14:  $\hat{\alpha}\leftarrow\alpha^{(i)}+h\,\tilde{v}^{\alpha}$
15:
16:  if i<S1i<S-1 then
17:   Corrector (Heun):
[…]
19:   $v^{\alpha}_{\mathrm{base},+}\leftarrow v^{\alpha}_{t_{i+1}}(\hat{x},\,\hat{\alpha},\,v^{x}_{+})$
20:   $(\tilde{v}^{x}_{+},\,\tilde{v}^{\alpha}_{+})\leftarrow v^{\text{ft}}_{t+1}(\hat{x},\,\hat{\alpha},\,v^{\alpha}_{\mathrm{base},+})$
21:   One-step terminal extrapolation:
22:   $\hat{\alpha}_{1}\leftarrow\hat{\alpha}+(1-t_{i+1})\,\tilde{v}^{\alpha}_{+}$
23:   Sparse-observation loss:
24:   $L\leftarrow\dfrac{1}{|\mathcal{I}|}\;\sum_{j\in\mathcal{I}}\bigl\|\alpha_{\mathrm{obs}}[j]-\hat{\alpha}_{1}[j]\bigr\|_{2}^{2}$
25:   Heun average update:
26:   $x^{(i+1)}\leftarrow x^{(i)}+\dfrac{h}{2}\,\bigl(\tilde{v}^{x}+\tilde{v}^{x}_{+}\bigr)$
27:   $\alpha^{(i+1)}\leftarrow\alpha^{(i)}+\dfrac{h}{2}\,\bigl(\tilde{v}^{\alpha}+\tilde{v}^{\alpha}_{+}\bigr)$
28:   Decaying guidance:
29:   $s\leftarrow\sqrt{\,1-i/S\,}$
30:   $x^{(i+1)}\leftarrow x^{(i+1)}-s\,\gamma_{x}\,\nabla_{x^{(i)}}L$
31:   $\alpha^{(i+1)}\leftarrow\alpha^{(i+1)}-s\,\gamma_{\alpha}\,\nabla_{\alpha^{(i)}}L$
32:  else
33:   Last step (no correction/guidance):
34:   $x^{(i+1)}\leftarrow\hat{x}$
35:   $\alpha^{(i+1)}\leftarrow\hat{\alpha}$
36:  end if
37:end for
38:return $\{x^{(i)}\}_{i=0}^{S}$, $\{\alpha^{(i)}\}_{i=0}^{S}$
E.5 Elasticity
Fine-tuning in the elasticity experiment is more challenging than the Darcy denoising experiment, which is why we increase the number of fine-tuning epochs to 100. Other parameters are the same as in the Darcy experiments.
Table 8: Elasticity fine-tuning hyperparameters.
Hyperparameter Value
Time–tilting factor $q$ 0.9
$\lambda_{x}$ varying
$\lambda_{\alpha}$ varying
$\lambda_{f}$ varying
$K_{\text{last}}$ 20
$K$ 20
Noise scaling $\kappa$ 0.9
Sampling steps (per trajectory) 100
Training epochs 100
Learning rate 0.00002
Batch size 15
We compare our fine-tuning approach with the inference-time correction method ECI presented in Cheng et al. (2024). Our implementation of the sampling correction method is given in Algorithm 3. For ECI, we set $M=5$ and use 1000 steps when sampling, compared to 100 steps used when sampling from the fine-tuned model. The reported residual heat maps in the main paper show one test function per grid point, where the magnitude indicates local error without aggregating across test functions.
Algorithm 3 ECI-style sampling with boundary correction
1:initial state $x_{0}$; steps $S$; inner correction iterations $M$; model drift $v_{t}(\cdot)$; correction operator $\mathcal{C}(\cdot)$; noise sampler $\operatorname{Noise}(\cdot)$
2:trajectory $\{x^{(i)}\}_{i=0}^{S}$
[…]
5:
6:for $i=0$ to S1S-1 do
7:  $t\leftarrow t_{i}$
8:  $t^{\text{next}}\leftarrow t_{i+1}$
9:
10:  Inner ECI corrections at fixed $t$:
11:  $\tilde{x}\leftarrow x^{(i)}$
12:  for $j=1$ to $M$ do
[…]
14:   $x_{\text{os}}\leftarrow\tilde{x}+(1-t)\,v$ one-step Euler update
15:   $x_{\text{corr}}\leftarrow\mathcal{C}\!\left(x_{\text{os}}\right)$ apply boundary correction
16:   $\eta\leftarrow\operatorname{Noise}(\text{shape of }x_{0})$
17:   $\tilde{x}\leftarrow(1-t)\,\eta+t\,x_{\text{corr}}$ stochastic convex blending
18:  end for
19:
20:  Final correction and roll-forward to $t^{\text{next}}$:
[…]
22:  $x_{\text{os}}\leftarrow\tilde{x}+(1-t)\,v$
23:  $x_{\text{corr}}\leftarrow\mathcal{C}\!\left(x_{\text{os}}\right)$
24:  $\eta\leftarrow\operatorname{Noise}(\text{shape of }x_{0})$
25:  $x^{(i+1)}\leftarrow(1-t^{\text{next}})\,\eta+t^{\text{next}}\,x_{\text{corr}}$
26:end for
27:return $\{x^{(i)}\}_{i=0}^{S}$
E.6 Helmholtz and Stokes Lid-driven Cavity
For both Helmholtz and the Stokes lid-driven cavity problems, we use the same configuration as in the elasticity experiment (see 8), training for 100 epochs. We only adjust the number of input/output channels.
E.7 Natural Images: Recoloring
For natural-image experiments, we adopt the class-conditional Latent Flow Matching (LFM) model of Dao et al. (2023) trained on ImageNet-1k (Deng et al., 2009) with a DiT backbone (Peebles and Xie, 2023). We fix an ImageNet class label $y$ to condition the generator and hold a global text prompt $c$ to define the fine-tuning direction. As a reward we use PickScore v1, a CLIP-H/14–based preference scorer trained on the Pick-a-Pic dataset, that evaluates image–text compatibility via cosine similarity in the CLIP embedding space (Kirstain et al., 2023; Radford et al., 2021; Cherti et al., 2023; Schuhmann et al., 2022), implemented with Transformers (Wolf et al., 2020).
Concretely, we maximize $R\Bigl(T\Bigl(D(z),\ \alpha\Bigr),\ c\Bigr)$, where D is the latent-space decoder and $T(\cdot,\alpha)$ is a parametric per-pixel recoloring with coefficients $\alpha$. The recoloring operates in logit space to avoid saturation: with $D(z)=x:\Omega\!\to\!(0,1)^{3}$,
$x_{\varepsilon}(\xi)=\mathrm{clip}\bigl(x(\xi),\varepsilon,1-\varepsilon\bigr),\qquad\ell(\xi)=\log\!\frac{x_{\varepsilon}(\xi)}{1-x_{\varepsilon}(\xi)}\in\mathbb{R}^{3},$
we apply a residual $r(\xi;\alpha)=\alpha\,\Phi_{D}^{\mathrm{bias}}\!\bigl(x(\xi)\bigr)$ built from RGB monomials up to total degree $D$ and map back via $y(\xi;\alpha)=\sigma\!\bigl(\ell(\xi)+r(\xi;\alpha)\bigr)$. Equivalently, for channel $c\in\{R,G,B\}$,
$y_{c}(\xi;\alpha)=\sigma\!\Bigl(\ell_{c}(\xi)+\sum_{m=1}^{M}\alpha_{cm}\,\phi_{m}(x(\xi))+\alpha_{c,0}\Bigr),\quad\phi_{m}\in\Phi_{D}.$
This residual parameterization is identity at initialization ($\alpha=0$) and provides a low-dimensional, channel-coupled appearance pathway that can disentangle content from presentation or reach colorings underrepresented by the base LFM. It is related to CNN-predicted polynomial color transforms (quadratic in Chai et al. (2020)), whereas we use cubic polynomials and learn solely via the reward.
The parameter predictor $\varphi$ here maps from latent feature tensors $z$ to parameters $\alpha$ . First, a compact scene descriptor is extracted by passing $z$ through convolutional transformations, aggregating information across multiple spatial scales via downsampling and global pooling, and enriching it with low-order channel statistics of $z$ (e.g., moments). The concatenated descriptor is projected into a fixed embedding space, refined by a lightweight pre–LayerNorm MLP with a residual connection, and finally mapped by a linear head to the recoloring coefficients $\alpha$.
Building on the inverse predictor, we augment the base DiT backbone with two lightweight residual heads that couple image dynamics and color–parameter evolution. First, the image path predicts the base drift $v_{t,x}^{\text{base}}(x,t)$. Then, we form a compact $\alpha$-token by flattening the current color parameters $\alpha$ and mapping them through a small MLP. This token is broadcast to a $k_{\alpha}$-channel feature map and concatenated with $(x,\,v_{t,x}^{\text{base}})$ into a shallow U–Net “correction” that outputs an additive refinement, yielding
$v_{t,x}^{\text{ft}}\;=\;v_{t,x}^{\text{base}}\;+\;\mathcal{U}_{x}\!\bigl(x,\,v_{t,x}^{\text{base}},\,\text{map}(\alpha),\,t\bigr).$
Second, the parameter path updates the polynomial recoloring coefficients by a residual on top of the baseline parameter field $v_{t,\alpha}^{\text{base}}$. Here, a dedicated $\alpha$-projection module mirrors the inverse predictor: multi-scale pooled conv features are fused—at the token level—with tokens from $\alpha$ and $v_{t,\alpha}^{\text{base}}$ via a small fusion MLP, AdaLN/FiLM modulation of a head token, and a light SE rescaling of conv features. The head then predicts a delta added to the baseline,
$v_{t,\alpha}^{\text{ft}}\;=\;v_{t,\alpha}^{\text{base}}\;+\;\mathcal{U}_{\alpha}\!\bigl(x,\,\alpha,\,v_{t,\alpha}^{\text{base}}\bigr).$
All correction heads are zero-initialized at their final projections, so fine-tuning starts from the base behavior and departs smoothly as training progresses. Table 9 lists the parameters of the correction U–Nets. In total, the modified architecture adds $\approx$9M parameters to the $\approx$130M parameters of the base DiT backbone.
Table 9: Correction head hyperparameters (image and parameter paths).
Hyperparameter Value
$k_{\alpha}$ 16
U–Net base widths [96, 128, 160, 192]
U–Net embedding 96
U–Net hidden lift 256
Attention stages [2, 3]
Attention heads 8
With the LFM model, we use 40 steps when sampling and use the same training specifications as the original Adjoint Matching paper (Domingo-Enrich et al., 2025). Again, we first pretrain the predictor $\varphi$, and then perform joint fine-tuning with our Adjoint Matching formulation. Fine-tuning is performed for 100 epochs with a batch size of 15.
Appendix F Additional Results
In this section, we provide full experimental results and samples from the trained models.
F.1 Full Experimental Results
F.1.1 Darcy
Figure 7: Darcy dataset: scatter plots showing the relationship between PDE residuals (weak and strong) and distributional discrepancies (MMDx, MMDα) across all model variants and hyperparameter configurations. Each point corresponds to one configuration; lower values indicate better physics consistency or distributional fidelity.
Table 10: Darcy dataset: full residual and distributional metrics for all methods and hyperparameter settings
Model $\lambda_{x}$ $\lambda_{f}$ $R_{\text{weak}}$ $\downarrow$ $R_{\text{strong}}$ $\downarrow$ MMDx $\downarrow$ MMDα $\downarrow$
FM $3.56\times 10^{0}$ ($\pm$ 0.62) $7.11\times 10^{1}$ ($\pm$ 3.20) $0.005$ $0.011$
PBFM $1.69\times 10^{0}$ ($\pm$ 0.14) $1.64\times 10^{1}$ ($\pm$ 1.66) $0.545$ $0.330$
Base AM 1k $1.89\times 10^{0}$ ($\pm$ 0.40) $5.55\times 10^{1}$ ($\pm$ 2.04) $0.238$ $0.057$
Base AM 5k $1.84\times 10^{0}$ ($\pm$ 0.55) $4.22\times 10^{1}$ ($\pm$ 1.63) $0.362$ $0.094$
Base AM 10k $1.95\times 10^{0}$ ($\pm$ 0.81) $2.95\times 10^{1}$ ($\pm$ 1.14) $0.380$ $0.114$
Base AM 20k $2.26\times 10^{0}$ ($\pm$ 1.38) $2.34\times 10^{1}$ ($\pm$ 1.10) $0.405$ $0.125$
Base AM 100k $3.44\times 10^{0}$ ($\pm$ 2.98) $2.04\times 10^{1}$ ($\pm$ 1.05) $0.507$ $0.145$
Base AM 1M $3.68\times 10^{0}$ ($\pm$ 3.17) $2.03\times 10^{1}$ ($\pm$ 1.04) $0.508$ $0.145$
Base AM 100M $3.72\times 10^{0}$ ($\pm$ 3.05) $2.02\times 10^{1}$ ($\pm$ 1.03) $0.512$ $0.146$
Base AM + $\varphi$ 1k $2.29\times 10^{0}$ ($\pm$ 0.42) $5.32\times 10^{1}$ ($\pm$ 1.89) $0.255$ $0.061$
Base AM + $\varphi$ 5k $1.68\times 10^{0}$ ($\pm$ 0.51) $4.01\times 10^{1}$ ($\pm$ 1.46) $0.376$ $0.095$
Base AM + $\varphi$ 10k $1.85\times 10^{0}$ ($\pm$ 0.66) $3.2\times 10^{1}$ ($\pm$ 1.16) $0.394$ $0.107$
Base AM + $\varphi$ 20k $2.15\times 10^{0}$ ($\pm$ 1.18) $2.26\times 10^{1}$ ($\pm$ 1.00) $0.409$ $0.122$
Base AM + $\varphi$ 100k $3.36\times 10^{0}$ ($\pm$ 2.61) $1.98\times 10^{1}$ ($\pm$ 1.03) $0.481$ $0.143$
Base AM + $\varphi$ 1M $3.82\times 10^{0}$ ($\pm$ 3.27) $1.87\times 10^{1}$ ($\pm$ 0.99) $0.518$ $0.148$
Base AM + $\varphi$ 100M $3.58\times 10^{0}$ ($\pm$ 3.06) $1.9\times 10^{1}$ ($\pm$ 1.01) $0.519$ $0.150$
AM 1k 0.0 $1.93\times 10^{0}$ ($\pm$ 0.32) $5.96\times 10^{1}$ ($\pm$ 2.60) $0.024$ $0.008$
AM 10k 0.0 $1.11\times 10^{0}$ ($\pm$ 0.27) $5.23\times 10^{1}$ ($\pm$ 2.18) $0.063$ $0.012$
AM 20k 0.0 $9.15\times 10^{-1}$ ($\pm$ 2.29) $4.71\times 10^{1}$ ($\pm$ 2.13) $0.053$ $0.010$
AM 20k 10 $1.97\times 10^{0}$ ($\pm$ 0.38) $5.13\times 10^{1}$ ($\pm$ 2.15) $0.010$ $0.010$
AM 20k 1 $1.64\times 10^{0}$ ($\pm$ 0.34) $4.94\times 10^{1}$ ($\pm$ 2.20) $0.012$ $0.009$
AM 20k $10^{-1}$ $1.23\times 10^{0}$ ($\pm$ 0.27) $4.73\times 10^{1}$ ($\pm$ 2.21) $0.023$ $0.006$
AM 20k $10^{-2}$ $1.05\times 10^{0}$ ($\pm$ 0.24) $4.63\times 10^{1}$ ($\pm$ 2.08) $0.049$ $0.009$
AM 20k $10^{-5}$ $9.78\times 10^{-1}$ ($\pm$ 2.18) $4.74\times 10^{1}$ ($\pm$ 2.19) $0.034$ $0.008$
AM 100k 0.0 $1.15\times 10^{0}$ ($\pm$ 0.28) $4.9\times 10^{1}$ ($\pm$ 2.58) $0.006$ $0.006$
AM 1M 0.0 $1.26\times 10^{0}$ ($\pm$ 0.32) $5.08\times 10^{1}$ ($\pm$ 2.72) $0.006$ $0.008$
AM 10M 0.0 $1.54\times 10^{0}$ ($\pm$ 0.35) $5.09\times 10^{1}$ ($\pm$ 2.69) $0.008$ $0.007$
AM 100M 0.0 $1.32\times 10^{0}$ ($\pm$ 0.31) $5.23\times 10^{1}$ ($\pm$ 2.77) $0.013$ $0.009$
F.1.2 Elasticity
Figure 8: Fine-tuning towards boundary modification, comparing our approach with ECI. Samples for two random seeds shared across models. Top row: displacement fields. Bottom row: corresponding weak residual heatmaps.
Figure 9: Elasticity dataset: scatter plots showing the relationship between PDE residuals (weak and strong) and distributional discrepancies (MMDx, MMDα) across all model variants and hyperparameter configurations. Each point corresponds to one configuration; lower values indicate better physics consistency or distributional fidelity.
Table 11: Elasticity dataset: full residual and distributional metrics for all methods and hyperparameter settings
Model $\lambda_{x}$ $\lambda_{f}$ $R_{\text{weak}}$ $\downarrow$ $R_{\text{strong}}$ $\downarrow$ MMDx $\downarrow$ MMDα $\downarrow$
FM $1.59\times 10^{1}$ ($\pm$ 0.37) $1.83\times 10^{1}$ ($\pm$ 0.66) $0.238$ $0.049$
PBFM $6.32\times 10^{0}$ ($\pm$ 0.82) $4.22\times 10^{0}$ ($\pm$ 0.26) $0.918$ $0.544$
Base AM 1k $6.78\times 10^{0}$ ($\pm$ 1.09) $1.21\times 10^{1}$ ($\pm$ 0.34) $0.325$ $0.194$
Base AM 10k $4.62\times 10^{0}$ ($\pm$ 0.85) $4.89\times 10^{0}$ ($\pm$ 1.29) $0.336$ $0.297$
Base AM 50k $5.47\times 10^{0}$ ($\pm$ 0.85) $4.16\times 10^{0}$ ($\pm$ 0.87) $0.404$ $0.314$
Base AM 100k $6.18\times 10^{0}$ ($\pm$ 0.75) $4.0\times 10^{0}$ ($\pm$ 0.98) $0.396$ $0.288$
Base AM 500k $6.68\times 10^{0}$ ($\pm$ 0.79) $4.04\times 10^{0}$ ($\pm$ 0.76) $0.417$ $0.299$
Base AM 1M $6.08\times 10^{0}$ ($\pm$ 0.74) $3.76\times 10^{0}$ ($\pm$ 0.54) $0.446$ $0.288$
Base AM + $\varphi$ 1k $6.99\times 10^{0}$ ($\pm$ 0.99) $1.06\times 10^{1}$ ($\pm$ 0.30) $0.323$ $0.180$
Base AM + $\varphi$ 10k $4.58\times 10^{0}$ ($\pm$ 0.70) $4.91\times 10^{0}$ ($\pm$ 1.31) $0.339$ $0.298$
Base AM + $\varphi$ 50k $5.55\times 10^{0}$ ($\pm$ 0.72) $4.58\times 10^{0}$ ($\pm$ 0.94) $0.359$ $0.290$
Base AM + $\varphi$ 100k $5.65\times 10^{0}$ ($\pm$ 0.82) $4.16\times 10^{0}$ ($\pm$ 0.93) $0.375$ $0.288$
Base AM + $\varphi$ 500k $5.47\times 10^{0}$ ($\pm$ 0.72) $3.92\times 10^{0}$ ($\pm$ 0.77) $0.408$ $0.296$
Base AM + $\varphi$ 1M $6.39\times 10^{0}$ ($\pm$ 0.86) $4.09\times 10^{0}$ ($\pm$ 0.68) $0.469$ $0.286$
AM 5k 1 $5.57\times 10^{0}$ ($\pm$ 0.60) $4.34\times 10^{0}$ ($\pm$ 0.95) $0.510$ $0.248$
AM 20k 1k $8.33\times 10^{0}$ ($\pm$ 1.34) $4.78\times 10^{0}$ ($\pm$ 1.15) $0.096$ $0.060$
AM 30k 1k $8.44\times 10^{0}$ ($\pm$ 1.27) $4.48\times 10^{0}$ ($\pm$ 1.08) $0.091$ $0.060$
AM 50k 1k $7.85\times 10^{0}$ ($\pm$ 1.17) $4.4\times 10^{0}$ ($\pm$ 1.08) $0.108$ $0.064$
AM 100k 0.0 $3.04\times 10^{0}$ ($\pm$ 0.31) $5.06\times 10^{0}$ ($\pm$ 0.66) $0.144$ $0.253$
AM 100k 1k $6.78\times 10^{0}$ ($\pm$ 1.00) $4.43\times 10^{0}$ ($\pm$ 1.02) $0.091$ $0.074$
AM 500k 1k $6.15\times 10^{0}$ ($\pm$ 0.77) $3.79\times 10^{0}$ ($\pm$ 0.87) $0.148$ $0.121$
AM 1M 0.0 $3.93\times 10^{0}$ ($\pm$ 0.27) $3.49\times 10^{0}$ ($\pm$ 0.48) $0.204$ $0.297$
AM 1M 1k $5.14\times 10^{0}$ ($\pm$ 0.42) $2.85\times 10^{0}$ ($\pm$ 0.58) $0.197$ $0.203$
F.1.3 Helmholtz
Figure 10: Helmholtz dataset: scatter plots showing the relationship between PDE residuals (weak and strong) and distributional discrepancies (MMDx, MMDα) across all model variants and hyperparameter configurations. Each point corresponds to one configuration; lower values indicate better physics consistency or distributional fidelity.
Table 12: Helmholtz dataset: full residual and distributional metrics for all methods and hyperparameter settings
Model $\lambda_{x}$ $\lambda_{f}$ $R_{\text{weak}}$ $\downarrow$ $R_{\text{strong}}$ $\downarrow$ MMDx $\downarrow$ MMDα $\downarrow$
FM $1.5\times 10^{1}$ ($\pm$ 0.59) $2.55\times 10^{1}$ ($\pm$ 0.55) $0.177$ $0.033$
PBFM $8.33\times 10^{0}$ ($\pm$ 3.04) $1.22\times 10^{1}$ ($\pm$ 0.33) $0.090$ $0.031$
Base AM 5k $5.64\times 10^{0}$ ($\pm$ 2.09) $1.59\times 10^{1}$ ($\pm$ 0.33) $0.134$ $0.039$
Base AM 50k $4.9\times 10^{0}$ ($\pm$ 1.85) $1.34\times 10^{1}$ ($\pm$ 0.32) $0.147$ $0.041$
Base AM 100k $5.19\times 10^{0}$ ($\pm$ 2.01) $1.27\times 10^{1}$ ($\pm$ 0.32) $0.148$ $0.042$
Base AM 1M $4.95\times 10^{0}$ ($\pm$ 1.81) $1.17\times 10^{1}$ ($\pm$ 0.33) $0.144$ $0.042$
Base AM 100M $5.01\times 10^{0}$ ($\pm$ 2.00) $1.14\times 10^{1}$ ($\pm$ 0.32) $0.143$ $0.042$
Base AM + $\varphi$ 5k $5.46\times 10^{0}$ ($\pm$ 1.94) $1.59\times 10^{1}$ ($\pm$ 0.33) $0.124$ $0.039$
Base AM + $\varphi$ 50k $5.35\times 10^{0}$ ($\pm$ 2.31) $1.35\times 10^{1}$ ($\pm$ 0.34) $0.133$ $0.045$
Base AM + $\varphi$ 100k $5.02\times 10^{0}$ ($\pm$ 2.17) $1.32\times 10^{1}$ ($\pm$ 0.33) $0.126$ $0.043$
Base AM + $\varphi$ 1M $4.99\times 10^{0}$ ($\pm$ 2.12) $1.16\times 10^{1}$ ($\pm$ 0.33) $0.125$ $0.046$
Base AM + $\varphi$ 100M $5.41\times 10^{0}$ ($\pm$ 2.20) $1.17\times 10^{1}$ ($\pm$ 0.34) $0.127$ $0.044$
AM 5k 1 $5.32\times 10^{0}$ ($\pm$ 1.49) $1.48\times 10^{1}$ ($\pm$ 0.29) $0.081$ $0.035$
AM 10k 0.0 $4.71\times 10^{0}$ ($\pm$ 1.26) $1.35\times 10^{1}$ ($\pm$ 0.28) $0.069$ $0.037$
AM 100k 0.0 $4.3\times 10^{0}$ ($\pm$ 1.29) $1.14\times 10^{1}$ ($\pm$ 0.29) $0.069$ $0.037$
AM 100k 0p1 $4.49\times 10^{0}$ ($\pm$ 1.46) $1.17\times 10^{1}$ ($\pm$ 0.29) $0.069$ $0.040$
AM 100k 1 $4.42\times 10^{0}$ ($\pm$ 1.35) $1.21\times 10^{1}$ ($\pm$ 0.30) $0.068$ $0.041$
AM 100k 1k $8.36\times 10^{0}$ ($\pm$ 2.65) $1.16\times 10^{1}$ ($\pm$ 0.25) $0.151$ $0.032$
AM 500k 0.0 $4.56\times 10^{0}$ ($\pm$ 1.47) $1.09\times 10^{1}$ ($\pm$ 0.31) $0.066$ $0.046$
AM 1M 0.0 $4.43\times 10^{0}$ ($\pm$ 1.43) $1.1\times 10^{1}$ ($\pm$ 0.30) $0.064$ $0.046$
AM 1M 1k $5.83\times 10^{0}$ ($\pm$ 1.63) $1.1\times 10^{1}$ ($\pm$ 0.27) $0.106$ $0.032$
AM 100M 0.0 $4.43\times 10^{0}$ ($\pm$ 1.43) $1.1\times 10^{1}$ ($\pm$ 0.30) $0.065$ $0.045$
AM 100M 1 $4.32\times 10^{0}$ ($\pm$ 1.43) $1.05\times 10^{1}$ ($\pm$ 0.30) $0.062$ $0.042$
F.2 Stokes Lid-driven Cavity
Figure 11: Stokes dataset: scatter plots showing the relationship between PDE residuals (weak and strong) and distributional discrepancies (MMDx, MMDα) across all model variants and hyperparameter configurations. Each point corresponds to a single configuration; lower values indicate better physics consistency or distributional fidelity. PBFM is omitted in the strong-residual panels because its training did not converge to physically meaningful samples (yielding strong residuals several orders of magnitude larger than the other methods.
[…]